# FiniteTimeBibliography.bib

### v1

@@ 1.1.1 Classical statistical mechanics without ergodicity
Statistical mechanics aims to reduce the behavior of complex systems to a simple effective description in terms of statistical ensembles. The question of how to establish the validity of such a statistical description in any given system remains of foundational interest. In classical statistical mechanics, the conventional justification — widely invoked in introductory accounts [Tolman, Reif, ReichlStatMech], though often with some hesitation [Tolman, LLStatMech] — is the ergodic hypothesis, originally due to Boltzmann. By supposing that a classical Hamiltonian system uniformly explores some region of phase space (often a surface of constant energy) over the course of its time evolution — in other words, shows ergodic dynamics — one attempts to justify describing the system by a uniform statistical distribution (the microcanonical distribution) over this region. If such a description is possible, the system is said to thermalize to this distribution.
While the ergodic hypothesis has led to a rich mathematical theory of ergodic classical dynamics [HalmosErgodic, SinaiCornfeld] in sufficiently simple systems [Sinai1976], it is an incredibly difficult problem to show ergodicity in systems with any realistic degree of complexity. Even if the ergodic hypothesis is assumed, the timescales required for the exploration of the phase space stretch far beyond the timescales at which statistical mechanics is known to be valid for typical observables of physical interest [KhinchinStatMech, GallavottiErgodic]. An alternate approach is to focus on a class of “physical” observables (determined according to what is feasible to measure in a given system), and examine when such observables may admit a statistical description — without any regard for whether the system itself shows ergodic dynamics [KhinchinStatMech, GallavottiErgodic]. For example, in a system of many particles, Khinchin [KhinchinStatMech, page 68] showed that if single-particle observables $a_{i}$ with respective mean values $\overline{a_{i}}$ (say) almost surely have vanishing “connected” autocorrelators at long times $t$:
@@
[…]
@@ 1.1.2 Eigenstate thermalization in quantum statistical mechanics
The conventional approach for describing the thermalization of Hamiltonian quantum systems is based on the eigenstate thermalization hypothesis (ETH) [vonNeumannThermalization, JensenShankarETH, deutsch1991eth, Deutsch2025Supplement, srednicki1994eth, srednicki1999eth, rigol2008eth, DAlessio2016, MoriETHreview, deutsch2018eth]. Here, it is important to refer to the eigenstates $\lvert E_{n}\rangle$ corresponding to the energy levels $E_{n}$ of the Hamiltonian $\hat{H}$. In the simplest version of ETH, formulated by von Neumann [vonNeumannThermalization], an observable $\hat{A}$ thermalizes to some thermal value $A_{\mathrm{th}}$ in every state in an infinite time average, provided that its expectation value thermalizes in every energy eigenstate:
@@
The more general statement of ETH [srednicki1999eth, rigol2008eth, DAlessio2016, MoriETHreview, deutsch2018eth] in wide present-day use, which is usually motivated [DAlessio2016] by comparison with random matrices — regarded as prototypes of complex quantum systems [Mehta] — posits the following structure of the matrix elements of $\hat{A}$ in the energy eigenbasis:
@@
Here, $A_{\mathrm{th}}(E)$ and $f_{A}(E_{1},E_{2})$ are smooth functions of energy, $R_{nm}$ is an appropriate standard random matrix with $O(1)$ matrix elements (at least to an initial approximation that neglects correlations), and $\exp(S(\overline{E}))$ is the (smooth) density of states around energy $\overline{E}=(E_{n}+E_{m})/2$, which is expected to scale with the Hilbert space dimension around this energy. It can be shown [srednicki1999eth] that if the energy levels $E_{n}$ have non-degenerate spacings, then Eq. (3) implies the thermalization of $\hat{A}$ to the thermal value $A_{\mathrm{th}}(E)$, at almost all times, for initial states supported near the energy $E$. Thus, in addition to eigenstate thermalization that is associated with infinite time averages, ETH also accounts for energy-dependence and thermal equilibrium. The hypothesis in ETH is that physically accessible observables in sufficiently complex systems satisfy Eq. (3).
@@
But which systems are we to count as “sufficiently” complex? Again, motivated by the idea of random matrices as prototypical complex systems, one approach is to assume that the appropriate systems are those whose energy level statistics resembles the eigenvalue statistics of random matrices [DAlessio2016]. However, even in such systems, it is straightforward to construct observables (such as projectors onto the energy eigenstates $\lvert E_{n}\rangle\langle E_{n}\rvert$) that do not satisfy Eq. (3), though these are expected to be generally inaccessible in experiments. The statistics of the energy levels themselves, being observable-independent, may instead be shown to be directly related to a form of ergodic dynamics in the Hilbert space [dynamicalqergodicity], reminiscent of the ergodic hypothesis. In particular, this notion is related to the ability to construct a Hermitian operator that approximately measures increments of time in the Hilbert space [^1] This is related to the intuitive notion of ergodicity as follows: in a classical system satisfying the ergodic hypothesis, phase space coordinates may be loosely used to measure the time that would have elapsed since the system started in the intial state.. Much like the ergodic hypothesis, we do not expect this property of dynamical quantum ergodicity to be easily accessible in sufficiently complex quantum systems, nor of immediate relevance to accessible few-particle observables. For example, measurements of energy level statistics have been proposed [SFFmeas, pSFF, AdwaySantos] and realized [pSFFexpt1] only in systems with a handful ($\lesssim$10$\lesssim$10$\lesssim$ 1010) of qubits. We also note that some of the aforementioned measurements [pSFF, AdwaySantos, pSFFexpt1] as well as fluctuation-dissipation relations [SchuckertETH] may yield indirect signatures of ETH, but not a conclusive experimental determination. In contrast, macroscopic properties of the energy levels unrelated to the ergodic hypothesis, corresponding to a finite energy resolution in experiments, are accessible [SFFmeas, pSFF] and set direct constraints on the dynamical emergence of statistical mechanics [dynamicalqspeedlimit, dynamicalqfastscrambling]. The takeaway lesson that we wish to emphasize from these observations is as follows: for questions of experimental relevance to quantum statistical mechanics, it is crucial to work with a finite energy resolution rather than the microscopic properties of individual energy levels.
@@
Separately, from a more theoretical perspective, we have argued in Ref. [dynamicalqergodicity] that (observable-independent) dynamical ergodicity and (observable-dependent) thermalization should be understood as fundamentally different notions for quantum systems, and their connection — if any — lies in additional, as yet unclear, physical restrictions on the class of interesting observables. Recent numerical studies suggest that natural restrictions such as locality are not sufficient to force a connection [MaganWu]. Moreover, one of the earliest numerical studies of eigenstate thermalization [JensenShankarETH] — which partly considers the question of how quantum thermalization may occur even in a system with accessible conserved quantities (intuitively “non-ergodic”) — observes eigenstate thermalization even in systems without the appropriate energy level distribution for dynamical ergodicity. For these reasons, although ergodicity remains an interesting dynamical property of quantum systems for fundamental reasons, such as for understanding nonperturbative quantum dynamics [bakeranomalieserg, JTreconstruction], we believe that an accessible approach to quantum thermalization should be independent of dynamical ergodicity (and therefore, the statistical properties of energy levels), and closer in spirit to Eq. (1).
@@
We will discuss questions of accessibility using the following rules of thumb, which are conventional for quantum many-body measurements [^2] We will also use the formal asymptotic notation [knuth1976asymptotic] where relevant. In the $x\to\infty$ limit, $y=o(x)$ represents $y<cx$ for any constant $c$, $y=O(x)$ represents $y\leq cx$ for some constant $c$, $y=\Theta(x)$ amounts to $c_{1}x\leq y\leq c_{2}x$ for constants $c_{1},c_{2}$, $y=\Omega(x)$ represents $y\geq cx$ for some $c$ and $y=\omega(x)$ represents $y>cx$ for any $c$.. For an $N$-particle system, we consider any quantity that requires a number of measurements, resolution or a time interval (relative to some experimentally natural timescale) that scales at most as $\operatorname{poly}(N)$ (representing some polynomial in $N$) to be accessible. The lower the degree of the polynomial, the more accessible the quantity. Somewhat more formally, if a quantity with value $Q\sim 1$ can be measured with a finite number of resources (which fixes the normalization of $Q$ that is relevant for questions of accessibility), then we expect that the following range of values of $Q$ are accessible:
@@
In contrast to the ergodic hypothesis, ETH in its very formulation addresses accessible observables. However, much like the ergodic hypothesis, it refers to the behavior of these observables in a large family of (typically) inaccessible states: the energy eigenstates and their simple superpositions (to determine the off-diagonal matrix elements). Correspondingly, ETH is also considered very hard to prove in generic cases, except in sufficiently simple models [MoriETHreview], especially in systems with certain special symmetry properties such as translation invariance [BiroliWeakETH, MoriWeakETH]. Even in many of these simpler models, the requirements of nondegeneracy of the energy levels and their spacings is a major obstacle for rigorously inferring thermalization from ETH: while nondegeneracy is believed to be generic [srednicki1999eth], it corresponds again to very fine properties of the energy levels that may be difficult to show definitively in any given system. Moreover, in an experiment, the resolution required to establish nondegeneracy is significantly higher than that required for the statistics of energy levels [pSFF]. Finally, even in the absence of degeneracies, as is believed to be the case for “typical” systems, the time after which ETH can guarantee thermalization tends to be inaccessibly large. Similar conclusions apply to other eigenstate-based approaches to thermalization [Borgonovi2016].
Let us now illustrate the magnitude of the problem with infinite time averages in ETH. Even assuming that somehow — despite their practical inaccessibility — we have a fairly good idea that an observable satisfies Eq. (3) and the system possesses nondegenerate energy level spacings, the standard constraint implied by ETH on the deviation of the expectation value $\langle\hat{A}(t)\rangle_{\rho}$ in an arbitrary state $\hat{\rho}$ from its thermal value $A_{\mathrm{th}}$ in the appropriate range of energies is (e.g., [DAlessio2016]):
@@
In other words, ETH cannot (rigorously) constrain thermal equilibrium until the same timescale $e^{S(\overline{E})}$ associated with dynamical ergodicity. Even in practice, one can roughly estimate [srednicki1999eth] the time scales of thermalization using the function $f(E_{1},E_{2})$, but only coupled with the properties of the initial state [DAlessio2016], which makes these estimates somewhat less “universal” than one may hope for. In terms of exact state-independent results, even for an observable that is known to satisfy the full statement of ETH, a guarantee of thermal equilibrium does not exist for accessible timescales due to Eq. (6). To our understanding, a similar shortcoming is present even in results on equilibration (without necessarily involving thermalization) [ReimannRealistic, ShortEqb] that rely on infinite time averages, though this can be mitigated to a certain extent with some knowledge of the spectrum and the initial state [ShortDegenerate].
@@ 1.2.1 Quantum thermalization with finite energy resolution
Figure 1: Energy-band thermalization depicted in terms of the relevant pairs of energy levels $(E_{n},E_{m})$ for two spectra with the same number of energy levels in the same range $[E_{\min},E_{\max}]$, but with (1(a)) having only occasional degeneracies, while (1(b)) is highly degenerate. The solid diagonal lines represent the energy band of interest around $E_{n}=E_{m}$ with a width $\Delta E$ [the orientation of the axes has been chosen to evoke matrix elements, which however would be a plot of $(n,m)$ instead of $(E_{n},E_{m})$; see also Fig. 5]. In general, it may not be possible to accessibly distinguish the two spectra (1(a)) and (1(b)) in an experiment, and the structure of energy level pairs inside the energy band can be completely different in each case. However, it is sufficient to know that whichever pairs of energy levels are in the band satisfy energy band thermalization, which is an experimentally accessible question, to conclusively determine thermalization in almost all initial states at all times larger than the inverse timescale of the energy band.
@@
Rather than connecting ETH to observable-independent ergodic dynamics or the statistics of energy levels as suggested by the traditional viewpoint [DAlessio2016], we will directly develop a connection between the experimentally accessible dynamics of quantum observables, similar to Eq. (1), and their thermalization in arbitrary physical basis states over experimentally accessible timescales. It is nevertheless convenient to express this structure in terms of the matrix elements of observables in the energy eigenstates (which form a preferred set of states uniquely identified by the dynamics of a system). For this purpose, we will introduce the notion of energy-band thermalization (whose structure is partly suggested by semiclassical proofs of eigenstate thermalization [ZelditchTransition, Sunada]) as a property of observables that does not directly depend on the energy levels of the system, but only on our (or an experimenter’s) choice of an energy band resolution $\Delta E$. We will also show that this determines thermalization over finite times $T\gg 2\pi/\Delta E$, without requiring infinite time averages (summarizing Sec. 4).
@@
Let us provide some context for why this should work. In the mathematical literature on semiclassical chaos [Shnirelman, CdV, ZelditchOG, ZelditchTransition, Sunada, ZelditchMixing, Zelditch, Anantharaman], it has been possible to prove eigenstate thermalization in the sense of Eq. (2) for classically accessible observables in certain quantum systems with a classical limit that satisfies the ergodic hypothesis, together with a (mild) suppression of off-diagonal elements in a shrinking energy band $\Delta E\to 0$ around the diagonal matrix elements. The collection of these two statements further implies classical ergodicity in the $T\to\infty$ limit. From our perspective, the fact that these statements successfully connect quantum matrix elements to purely classical dynamics strongly suggests [^3] Why this is suggested by classical theorems is roughly as follows: in these theorems, the classical limit $\hbar\to 0$, which is very similar to the thermodynamic limit $N\to\infty$ in that both increase the Hilbert space dimension to infinity, is taken before the $t\to\infty$ limit in these proofs. The classical infinite time average does not correspond to timescales that diverge as $\hbar\to 0$ (which are already set to $\infty$ when taking the classical limit), and in fact necessarily converges over $O(\hbar^{0})$ timescales, therefore being quite different from the quantum infinite time average. This means that the timescales referred to as “infinite” in the classical limit are actually well below the timescales needed to resolve the energy levels (which grow as a power of 1/$1/\hbar$ by the uncertainty principle for quantized classical systems, and as $\exp(N)$ for fully quantum many-body systems). that these are appropriate notions for describing quantum thermalization even when the quantization of the energy levels is inaccessible, as in the $\hbar\to 0$ limit. As we aim to describe finite dimensional quantum systems, we do not have an explicit analogue of Planck’s constant $\hbar$ available to us to safely take limits such as $T\to\infty$ without crossing the resolution of energy levels (we prefer not to take the thermodynamic limit of $N\to\infty$ particles in our results to retain applicability to finite but large systems as may be accessed in experiments); therefore, we must formulate our approach strictly for finite energy bands $\Delta E$ and finite time scales $T$.
@@
$[\hat{A}]_{\Delta E}\equiv\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\lvert E_{n}\rangle\langle E_{n}\rvert\hat{A}\lvert E_{m}\rangle\langle E_{m}\rvert.$ (7)
@@
$[\hat{A}]_{\Delta E}\approx A_{\mathrm{th}}\hat{\mathds{1}}.$ (8)
@@
$[\hat{A}\otimes\hat{A}]_{\Delta E}\approx A_{\mathrm{th}}^{2}\hat{\mathds{1}}\otimes\hat{\mathds{1}}.$ (10)
@@
Here, “almost all $t$” should be interpreted as “all $t$ except a set whose length is smaller than $T/c_{2}$ for some large constant $c_{2}>1$”.
This “cloning” strategy may be regarded as a quantum version of a theorem in classical ergodic theory [SinaiCornfeld], which states that a dynamical system is weakly mixing (which loosely corresponds to thermal equilibrium) if and only if its “cloned” version with two copies of the system is ergodic (which loosely corresponds to time-averaged thermalization). This is especially convenient for us, as it allows us to use the same conceptual notion of energy-band thermalization (but applied to the observable vs. its cloned version) to address both time-averaged thermalization and instantaneous thermal equilibrium over finite timescales.

### v2

# Bypassing eigenstate thermalization with experimentally accessible quantum dynamics

@@ 1.1.1 Classical statistical mechanics without ergodicity
Statistical mechanics aims to reduce the behavior of complex systems to a simple effective description in terms of statistical ensembles. The question of how to establish the validity of such a statistical description in any given system remains of foundational interest. In classical statistical mechanics, the conventional justification — widely invoked in introductory accounts [1, 2, 3], though often with some hesitation [1, 4] — is the ergodic hypothesis, originally due to Boltzmann. By supposing that a classical Hamiltonian system uniformly explores some region of phase space (often a surface of constant energy) over the course of its time evolution — in other words, shows ergodic dynamics — one attempts to justify describing the system by a uniform statistical distribution (the microcanonical distribution) over this region. If such a description is possible, the system is said to thermalize to this distribution.
While the ergodic hypothesis has led to a rich mathematical theory of ergodic classical dynamics [5, 6] in sufficiently simple systems [7], it is an incredibly difficult problem to show ergodicity in systems with any realistic degree of complexity. Even if the ergodic hypothesis is assumed, the timescales required for the exploration of the phase space stretch far beyond the timescales at which statistical mechanics is known to be valid for typical observables of physical interest [8, 9]. An alternate approach is to focus on a class of “physical” observables (determined according to what is feasible to measure in a given system), and examine when such observables may admit a statistical description — without any regard for whether the system itself shows ergodic dynamics [8, 9]. For example, in a system of many particles, Khinchin [8, page 68] showed that if single-particle observables $a_{i}$ with respective mean values $\overline{a_{i}}$ (say) almost surely have vanishing “connected” autocorrelators at long times $t$:
@@
where the average $\langle\cdot\rangle$ is over all trajectories in the phase space, then any observable constructed as $A=\sum_{i}c_{i}a_{i}$ may be described by the microcanonical ensemble over long times, in almost all many-particle initial states.
@@ 1.1.2 Eigenstate thermalization in quantum statistical mechanics
The conventional approach for describing the thermalization of Hamiltonian quantum systems is based on the eigenstate thermalization hypothesis (ETH) [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Here, it is important to refer to the eigenstates $\lvert E_{n}\rangle$ corresponding to the energy levels $E_{n}$ of the Hamiltonian $\hat{H}$. In the simplest version of ETH, formulated by von Neumann [10], an observable $\hat{A}$ thermalizes to some thermal value $A_{\mathrm{th}}$ in every state in an infinite time average, provided that its expectation value thermalizes in every energy eigenstate:
@@
The more general statement of ETH [15, 16, 17, 18, 19] in wide present-day use, which is usually motivated [17] by comparison with random matrices — regarded as prototypes of complex quantum systems [20] — posits the following structure of the matrix elements of $\hat{A}$ in the energy eigenbasis:
@@
Here, $A_{\mathrm{th}}(E)$ and $f_{A}(E_{1},E_{2})$ are smooth functions of energy, $R_{nm}$ is an appropriate standard random matrix with $O(1)$ matrix elements (at least to an initial approximation that neglects correlations), and $\exp(S(\overline{E}))$ is the (smooth) density of states around energy $\overline{E}=(E_{n}+E_{m})/2$, which is expected to scale with the Hilbert space dimension around this energy. It can be shown [15] that if the energy levels $E_{n}$ have non-degenerate spacings, then Eq. (3) implies the thermalization of $\hat{A}$ to the thermal value $A_{\mathrm{th}}(E)$, at almost all times, for initial states supported near the energy $E$. Thus, in addition to eigenstate thermalization that is associated with infinite time averages, ETH also accounts for energy-dependence and thermal equilibrium. The hypothesis in ETH is that physically accessible observables in sufficiently complex systems satisfy Eq. (3).
@@
But which systems are we to count as “sufficiently” complex? Again, motivated by the idea of random matrices as prototypical complex systems, one approach is to assume that the appropriate systems are those whose energy level statistics resembles the eigenvalue statistics of random matrices [17]. However, even in such systems, it is straightforward to construct observables (such as projectors onto the energy eigenstates $\lvert E_{n}\rangle\langle E_{n}\rvert$) that do not satisfy Eq. (3), though these are expected to be generally inaccessible in experiments. The statistics of the energy levels themselves, being observable-independent, may instead be shown to be directly related to a form of ergodic dynamics in the Hilbert space [21], reminiscent of the ergodic hypothesis. In particular, this notion is related to the ability to construct a Hermitian operator that approximately measures increments of time in the Hilbert space [^1] This is related to the intuitive notion of ergodicity as follows: in a classical system satisfying the ergodic hypothesis, phase space coordinates may be loosely used to measure the time that would have elapsed since the system started in the intial state.. Much like the ergodic hypothesis, we do not expect this property of dynamical quantum ergodicity to be easily accessible in sufficiently complex quantum systems that are thermodynamically large, nor of immediate relevance to accessible few-particle observables. For example, measurements of energy level statistics have been proposed [22, 23, 24] and realized [25] only in systems with a handful ($\lesssim$$\lesssim 10$) of qubits. We also note that some of the aforementioned measurements [23, 24, 25] as well as fluctuation-dissipation relations [26] may yield indirect signatures of ETH, but not a conclusive experimental determination. In contrast, macroscopic properties of the energy levels unrelated to the ergodic hypothesis, corresponding to a finite energy resolution in experiments, are accessible [22, 23] and set direct constraints on the dynamical emergence of statistical mechanics [27, 28]. The takeaway lesson that we wish to emphasize from these observations is as follows: for questions of experimental relevance to quantum statistical mechanics, it is crucial to work with a finite energy resolution rather than the microscopic properties of individual energy levels.
@@
Separately, from a more theoretical perspective, we have argued in Ref. [21] that (observable-independent) dynamical ergodicity and (observable-dependent) thermalization should be understood as fundamentally different notions for quantum systems, and their connection — if any — lies in additional, as yet unclear, physical restrictions on the class of interesting observables. Recent numerical studies suggest that natural restrictions such as locality are not sufficient to force a connection [29]. Moreover, one of the earliest numerical studies of eigenstate thermalization [11] — which partly considers the question of how quantum thermalization may occur even in a system with accessible conserved quantities (intuitively “non-ergodic”) — observes eigenstate thermalization even in systems without the appropriate energy level distribution for dynamical ergodicity. For these reasons, although ergodicity as determined by spectral statistics remains an interesting dynamical property of quantum systems for other fundamental reasons, such as for understanding nonperturbative quantum dynamics [30, 31] or for preparing highly entangled states for quantum teleportation in smaller complex systems [32], we believe that an accessible approach to quantum statistical mechanics should be independent of dynamical ergodicity (and therefore, the statistical properties of energy levels), and closer in spirit to Eq. (1).
@@
We will discuss questions of accessibility using the following rules of thumb, which are conventional for quantum many-body measurements [^2] We will also use the formal asymptotic notation [33] where relevant. In the $x\to\infty$ limit, $y=o(x)$ represents $y<cx$ for any constant $c$, $y=O(x)$ represents $y\leq cx$ for some constant $c$, $y=\Theta(x)$ amounts to $c_{1}x\leq y\leq c_{2}x$ for constants $c_{1},c_{2}$, $y=\Omega(x)$ represents $y\geq cx$ for some $c$ and $y=\omega(x)$ represents $y>cx$ for any $c$.. For an $N$-particle system, we consider any quantity that requires a number of measurements, resolution or a time interval (relative to some experimentally natural timescale) that scales at most as $\operatorname{poly}(N)$ (representing some polynomial in $N$) to be accessible. The lower the degree of the polynomial, the more accessible the quantity. Somewhat more formally, if a quantity with value $Q\sim 1$ can be measured with a finite number of resources (which fixes the normalization of $Q$ that is relevant for questions of accessibility), then we expect that the following range of values of $Q$ are accessible:
@@
In contrast to the ergodic hypothesis, ETH in its very formulation addresses accessible observables. However, much like the ergodic hypothesis, it refers to the behavior of these observables in a large family of (typically) inaccessible states: the energy eigenstates and their simple superpositions (to determine the off-diagonal matrix elements). Correspondingly, ETH is also considered very hard to prove in generic cases, except in sufficiently simple models [18], especially in systems with certain special symmetry properties such as translation invariance [34, 35]. Even in many of these simpler models, the requirements of nondegeneracy of the energy levels and their spacings is a major obstacle for rigorously inferring thermalization from ETH: while nondegeneracy is believed to be generic [15], it corresponds again to very fine properties of the energy levels that may be difficult to show definitively in any given system. Moreover, in an experiment, the resolution required to establish nondegeneracy is significantly higher than that required for the statistics of energy levels [23]. Finally, even in the absence of degeneracies, as is believed to be the case for “typical” systems, the time after which ETH can guarantee thermalization tends to be inaccessibly large. Similar conclusions apply to other eigenstate-based approaches to thermalization [36].
Let us now illustrate the magnitude of the problem with infinite time averages in ETH. Even assuming that somehow — despite their practical inaccessibility — we have a fairly good idea that an observable satisfies Eq. (3) and the system possesses nondegenerate energy level spacings, the standard constraint implied by ETH on the deviation of the expectation value $\langle\hat{A}(t)\rangle_{\rho}$ in an arbitrary state $\hat{\rho}$ from its thermal value $A_{\mathrm{th}}$ in the appropriate range of energies is (e.g., [17]):
@@
In other words, ETH cannot (rigorously) constrain thermal equilibrium until the same timescale $e^{S(\overline{E})}$ associated with dynamical ergodicity. Even in practice, one can roughly estimate [15] the time scales of thermalization using the function $f(E_{1},E_{2})$, but only coupled with the properties of the initial state [17], which makes these estimates somewhat less “universal” than one may hope for. In terms of exact state-independent results, even for an observable that is known to satisfy the full statement of ETH, a guarantee of thermal equilibrium does not exist for accessible timescales due to Eq. (6). To our understanding, a similar shortcoming is present even in results on equilibration (without necessarily involving thermalization) [37, 38] that rely on infinite time averages, though this can be mitigated to a certain extent with some knowledge of the spectrum and the initial state [39].
@@ 1.2.1 Quantum thermalization with finite energy resolution
Figure 1: Energy-band thermalization depicted in terms of the relevant pairs of energy levels $(E_{n},E_{m})$ for two spectra with the same number of energy levels in the same range $E_{\min},E_{\max}]$, but with (1(a)) having only occasional degeneracies, while (1(b)) is highly degenerate. The solid diagonal lines represent the energy band of interest around $E_{n}=E_{m}$ with a width $\Delta E$ [the orientation of the axes has been chosen to evoke matrix elements, which however would be a plot of $(n,m)$ instead of $(E_{n},E_{m})$; see also Fig. 5]. In general, it may not be possible to accessibly distinguish the two spectra (1(a)) and (1(b)) in an experiment, and the structure of energy level pairs inside the energy band can be completely different in each case. However, it is sufficient to know that whichever pairs of energy levels are in the band satisfy energy band thermalization, which is an experimentally accessible question, to conclusively determine thermalization in almost all initial states at all times larger than the timescale corresponding to the inverse width of the energy band.
@@
Rather than connecting ETH to observable-independent ergodic dynamics or the statistics of energy levels as suggested by the traditional viewpoint [17], we will directly develop a connection between the experimentally accessible dynamics of quantum observables, similar to Eq. (1), and their thermalization in arbitrary physical basis states over experimentally accessible timescales. It is nevertheless convenient to express this structure in terms of the matrix elements of observables in the energy eigenstates (which form a preferred set of states uniquely identified by the dynamics of a system). For this purpose, we will introduce the notion of energy-band thermalization (whose structure is partly suggested by semiclassical proofs of eigenstate thermalization [40, 41]) as a property of observables that does not directly depend on the energy levels of the system, but only on our (or an experimenter’s) choice of an energy band resolution $\Delta E$. We will also show that this determines thermalization over finite times $T\gg 2\pi/\Delta E$, without requiring infinite time averages (summarizing Sec. 4).
@@
Let us provide some context for why this should work. In the mathematical literature on semiclassical chaos [42, 43, 44, 40, 41, 45, 46, 47], it has been possible to prove eigenstate thermalization in the sense of Eq. (2) for classically accessible observables in certain quantum systems with a classical limit that satisfies the ergodic hypothesis, together with a (mild) suppression of off-diagonal elements in a shrinking energy band $\Delta E\to 0$ around the diagonal matrix elements. The collection of these two statements further implies classical ergodicity in the $T\to\infty$ limit. From our perspective, the fact that these statements successfully connect quantum matrix elements to purely classical dynamics strongly suggests [^3] Why this is suggested by classical theorems is roughly as follows: in these theorems, the classical limit $\hbar\to 0$, which is very similar to the thermodynamic limit $N\to\infty$ in that both increase the Hilbert space dimension to infinity, is taken before the $t\to\infty$ limit in these proofs. The classical infinite time average does not correspond to timescales that diverge as $\hbar\to 0$ (which are already set to $\infty$ when taking the classical limit), and in fact necessarily converges over $O(\hbar^{0})$ timescales, therefore being quite different from the quantum infinite time average. This means that the timescales referred to as “infinite” in the classical limit are actually well below the timescales needed to resolve the energy levels (which grow as a power of $1/\hbar$ by the uncertainty principle for quantized classical systems, and as $\exp(N)$ for fully quantum many-body systems). that these are appropriate notions for describing quantum thermalization even when the quantization of the energy levels is inaccessible, as in the $\hbar\to 0$ limit. As we aim to describe finite dimensional quantum systems, we do not have an explicit analogue of Planck’s constant $\hbar$ available to us to safely take limits such as $T\to\infty$ without crossing the resolution of energy levels (we prefer not to take the thermodynamic limit of $N\to\infty$ particles in our results to retain applicability to finite but large systems as may be accessed in experiments); therefore, we must formulate our approach strictly for finite energy bands $\Delta E$ and finite time scales $T$.
@@
$\hat{A}]_{\Delta E}\equiv\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\lvert E_{n}\rangle\langle E_{n}\rvert\hat{A}\lvert E_{m}\rangle\langle E_{m}\rvert$. (7)
@@
$\hat{A}]_{\Delta E}\approx A_{\mathrm{th}}\hat{\mathds{1}}$. (8)
@@
$\hat{A}\otimes\hat{A}]_{\Delta E}\approx A_{\mathrm{th}}^{2}\hat{\mathds{1}}\otimes\hat{\mathds{1}}$. (10)
@@
Here, “almost all tt” should be interpreted as “all $t$ except a set whose length is smaller than $T/c_{2}$ for some large constant $c_{2}>1$”.
This “cloning” strategy may be regarded as a quantum version of a theorem in classical ergodic theory [6], which states that a dynamical system is weakly mixing (which loosely corresponds to thermal equilibrium) if and only if its “cloned” version with two copies of the system is ergodic (which loosely corresponds to time-averaged thermalization). This is especially convenient for us, as it allows us to use the same conceptual notion of energy-band thermalization (but applied to the observable vs. its cloned version) to address both time-averaged thermalization and instantaneous thermal equilibrium over finite timescales.